# Fast Bilateral Teleoperation and Imitation Learning Using Sensorless Force Control via Accurate Dynamics Model

### v2

@@ 1 Introduction
However, most of these systems used unilateral control, which only transmits target position values from the leader robot to the follower robot [9]. This method is easy to implement and provides good operability in low-speed, low-contact operations, but it does not provide force feedback and struggles with contact-rich tasks. In fact, it also has difficulties in high-speed operations. On the other hand, there has been an increase in cases where bilateral control, which enables two-way information exchange between the leader robot and the follower robot and provides force feedback, has been adopted. There are various types of bilateral control. In some cases, force reflecting type bilateral control is used, in which external forces from followers are generated feedforward-style on the leader side, like FACTR [10] . However, this method is known to have stability issues, and in many cases, stability is ensured by adding a damping term [10][11], but this tends to impair operability. On the other hand, 4-channel bilateral control is a method in which both the leader and follower transmit their respective positions and forces to each other and perform position control and force control in parallel. Theoretically, a 4-channel architecture can achieve the ideal response of synchronization of position and force perfectly [12][13]. However, to perform position control and force control with opposite objectives simultaneously in a single robot, making them non-interfering requires accurate dynamics models and observation of position, velocity, and external forces. Because the dynamics of manipulators with six or more degrees of freedom involve many parameters and extremely complex functions, obtaining accurate parameters and performing real-time calculations has long been impractical. As a practical implementation method, an accelerated control system was proposed that combines a simplified dynamics model based on linear models of each joint with disturbance estimation (including external force and model error) using a disturbance observer [14] and compensation using the estimated disturbance. This enables high performance even when a certain amount of model error exists [15][16][17][18][19]. This acceleration control-based 4-channel bilateral control system has also been studied for application to imitation learning. Sakaino et al. trained a Long-Short-Term Memory (LSTM) [20] using data collected with four-channel bilateral control and conducted experiments to perform contact-rich tasks [21][22][23]. Kobayashi et al. published Alpha-$\alpha$ and Bi-ACT series [24][25][26][27] that are a combination of a low-cost 4-channel bilateral control system Alpha-$\alpha$ and the method to train Action Chunking with Transformer (ACT) [5] using data collected by 4-channel bilateral control.
@@ 2 Methodology
where $\dot{\bm{\theta}}_{-}$ and $\dot{\bm{\theta}}_{+}$ are difference and average of the joint angler velocity. This transformation allows us to think of position and force control independently. In other words, if the same amount of velocity is applied to both the leader and the follower in opposite directions, only the difference in the joint angle will change, not the average. Conversely, if the same amount of velocity is applied to both the leader and the follower in the same direction, only the average joint angle will change, not the difference. Ideally, these do not influence each other, i.e., they are orthogonal, so control systems can be designed separately for each. Since the coordinate transformation matrix is time-invariant, the coordinate transformation of the angle and angular acceleration can be described by the same matrix $\bm{J}_{+-}$. The dynamics of the difference and the average of the joint angles are as follows.
@@
$\begin{bmatrix}\bm{\tau}_{-}\\ \bm{\tau}_{+}\\ \end{bmatrix}_{ext}=\bm{J}_{+-}^{-\top}\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}=\begin{bmatrix}\frac{1}{2}\bm{I}&-\frac{1}{2}\bm{I}\\ \bm{I}&\bm{I}\\ \end{bmatrix}\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}=\begin{bmatrix}\bm{M}_{-}&\bm{0}\\ \bm{0}&\bm{M}_{+}\\ \end{bmatrix}_{des}\begin{bmatrix}\ddot{\bm{\theta}}_{-}\\ \ddot{\bm{\theta}}_{+}\\ \end{bmatrix}+\begin{bmatrix}\bm{D}_{-}&\bm{0}\\ \bm{0}&\bm{0}\\ \end{bmatrix}_{des}\begin{bmatrix}\dot{\bm{\theta}}_{-}\\ \dot{\bm{\theta}}_{+}\\ \end{bmatrix}+\begin{bmatrix}\bm{K}_{-}&\bm{0}\\ \bm{0}&\bm{0}\\ \end{bmatrix}_{des}\begin{bmatrix}\bm{\theta}_{-}\\ \bm{\theta}_{+}\\ \end{bmatrix}$ (6)
where $\bigcirc_{des}$ are desired dynamics parameters, and $\bm{\tau}_{-,ext},\bm{\tau}_{+,ext}$ are external forces corresponding to the difference and average of the joint angle. Since the system has position control and force control on orthogonal axes, the system can also be interpreted as position/force hybrid control [31] on transformed coordinates [17]. The control law can be derived by substituting the desired dynamics into the transformed dynamics:
$\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{u}$ $=\begin{bmatrix}\bm{M}_{l}(\bm{\theta}_{l})&\bm{0}\\ \bm{0}&\bm{M}_{f}(\bm{\theta}_{f})\\ \end{bmatrix}\begin{bmatrix}\frac{1}{2}\bm{I}&\bm{I}\\ -\frac{1}{2}\bm{I}&\bm{I}\\ \end{bmatrix}\begin{bmatrix}-\bm{M}_{-,des}^{-1}\bm{D}_{-,des}\dot{\bm{\theta}}_{-}-\bm{M}_{-,des}^{-1}\bm{K}_{-,des}\bm{\theta}_{-}\\ \bm{M}_{+,des}^{-1}\bm{\tau}_{+}\\ \end{bmatrix}-\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}$ (7)
@@
where $\bm{K}_{p}=\frac{1}{2}\bm{M}_{-,des}^{-1}\bm{K}_{-,des}$, $\bm{K}_{d}=\frac{1}{2}\bm{M}_{-,des}^{-1}\bm{D}_{-,des}$, and $\bm{K}_{f}=\bm{M}_{+,des}^{-1}$. Finally, the control law for each manipulator can be as follows:
$\bm{\tau}_{ref}$ $=\tilde{\bm{M}}(\bm{\theta})\left\{\bm{K}_{p}(\bm{\theta}_{des}-\bm{\theta})+\bm{K}_{d}(\dot{\bm{\theta}}_{des}-\hat{\dot{\bm{\theta}}})+\bm{K}_{f}(\bm{\tau}_{des}+\hat{\bm{\tau}}_{ext})\right\}-\hat{\bm{\tau}}_{ext}+\tilde{\bm{C}}(\bm{\theta},\hat{\dot{\bm{\theta}}})\hat{\dot{\bm{\theta}}}+\tilde{\bm{D}}\hat{\dot{\bm{\theta}}}+\tilde{\bm{g}}(\bm{\theta})$ (9)
@@ 2.2 Velocity and External Force Observer
Here, the observer gain is set using $\omega_{c}$, corresponding to the cut-off angular frequency in the frequency domain. The critical point is that inertia is not a constant but a function of the joint angle vector. Conventional disturbance observers use constants for inertia, and the effects of modeling errors are included in the estimated external force. Therefore, it was assumed that inertia could be nominalized by feedback compensation of the external force. However, the extent to which modeling errors can be compensated depends on how high the cut-off frequency (observer gain) can be set. In practice, when using control equipment capable of achieving high-precision rotary encoders and high sampling frequencies, it is possible to set a fairly high cut-off frequency, and by combining fast external force estimation and compensation for modeling errors, it is practically feasible to ignore modeling errors. However, in the case of low-cost hardware like the one used here, a high cut-off frequency cannot be set, so the error in inertia has a significant impact that cannot be ignored. By pre-identifying the nonlinear dynamics of the manipulator with 38 parameters, including the inertial matrix, and using this to always employ accurate inertial parameters, it is possible to improve control performance while maintaining a low cost.
Furthermore, using the identified inertia justifies the use of a velocity observer. In many studies of 4-channel bilateral control, pseudo-differentiation ($\frac{\omega_{c}s}{s+\omega_{c}}$) is used for velocity estimation, but pseudo-differentiation is known to cause estimation delays [32]. Nevertheless, pseudo-differentiation is widely used because the inertia is not accurately known, and methods that can calculate using only angle response values without using inertia are preferred. If accurate inertia can be obtained, it becomes possible to use not only the information from the angular response values but also predictions from the torque reference values, thereby improving the delay in speed estimation. Improving the delay in speed estimation allows the D gain of the position PD control to be set higher, and setting the D gain higher enables the P gain to be set higher as well, leading to a significant improvement in position control performance.
Converting this to a transfer function expression using the Laplace transform, we can write it as follows.
@@
We implemented the algorithm using a bilinear transformation ($s=\frac{2}{T}\cdot\frac{1-z^{-1}}{1+z^{-1}}$).
@@ 3 Experiments
Figure 3: Results of teleoperation comparison
@@ 3.1 Teleoperation Performance
Table 1: Success rate of dual arm pick-and-place tasks
Action Chunking with Transformer
@@
pick place pick place pick place
10mm 0/5 0/5 5/5 0/5 5/5 5/5
20mm 0/5 0/5 5/5 5/5 5/5 5/5
30mm 5/5 0/5 5/5 5/5 5/5 5/5
40mm 5/5 0/5 5/5 4/5 5/5 5/5
50mm 5/5 0/5 5/5 0/5 5/5 5/5
@@
Table 2: Success rate of imitation learning tasks
ACT
[…]
@@
3.2 Imitation Learning
We experimented to collect demonstration data using the proposed 4-channel bilateral teleoperation method and perform imitation learning. We used Action Chunking with Transformer (ACT) [5] for neural network architecture. The combination of ACT and 4-channel bilateral control is called Bi-ACT [24], and has already been validated on Pick-and-Place and Put-in-drawer tasks. With 4-channel bilateral control, data on the position and force of the leader and follower can be obtained. In addition to using all of this data, we also performed learning using only position information and force information only on the input side. We investigated the effects of using force information on both the input and output sides.
It is worth noting that a control system utilizing position and force command values was employed during data collection. Since the control system differs between data collection and autonomous operation, a decrease in success rate was expected when force information is not used in the policy output. However, differences in control systems or dynamics between data collection and autonomous operation are common when using data collection methods other than teleoperation, and this is not uncommon in imitation learning. Additionally, since the dataset’s quality can significantly affect imitation learning results, it is preferable to compare using the same dataset. Considering these points, we conducted this comparison.
Task1: Dual-arm Pick-and-Place. We performed a task of grasping blocks of different widths between two arms. We prepared five types of blocks, ranging from 10 mm to 50 mm in width, and collected a total of 10 demonstration data sets, two for each type. The blocks are shown in Fig. 8, and the results are shown in Table 2. As a result, when no force information was used for either input or output, picking was not possible in cases with small widths of 10 mm and 20 mm. In cases with relatively large widths of 30 mm to 50 mm, picking was possible once, but problems such as dropping the object immediately or the angle changing while holding it occurred, and the object could not be placed in the correct position. When force information was used in the input, the success rate of picking improved significantly, with all trials resulting in successful picking. However, instability after picking persisted, and placement failures occurred frequently in cases where the width was farthest from the average of 10mm and 50mm. When both input and output used force information, both pick and place were successful in all trials. The stability of holding after picking improved, with no instances of dropping midway or misalignment.
Task2: Nut Turning. We tried the nut turning task as a task that requires more force and speed. The results are shown in the first row of Table 2. As a result, when no force was applied to either the input or output, the experiment was completely unsuccessful, but when force was applied to the input, the experiment was successful regardless of whether force was applied to the output. When no force was applied to the input, the hand often stopped at the position where it overlapped with the nut on the bird’s-eye view image, and it seemed difficult to touch the nut due to insufficient force information.
Task3: Cucumber Peeling. As a task involving handling irregularly shaped objects, we tried peeling cucumbers with a peeler. The results are shown in the third row of Table 2. As a result, when no force information was used for either input or output, the experiment was completely unsuccessful. When force information was used for input, the experiment was successful in 3/5 cases, and when force information was used for both input and output, the experiment was successful in 2/5 cases. The main failure case was when the system transitioned to the finished state before touching the cucumber or peeler. This is thought to be because there was not much difference between the image at the start of the task and the image after the task was completed.
@@ 4 Conclusion
In this study, we demonstrated that a low-cost manipulator can achieve high-speed teleoperation with force feedback without force sensors by incorporating a high-precision manipulator dynamics model into a four-channel bilateral teleoperation system. We compared the proposed method to three teleoperation methods, such as unilateral control, symmetric position type, and force feedback type, with a 4-channel system that does not utilize manipulator dynamics. We confirmed that the proposed method demonstrates higher position and force tracking performance during high-speed operations. Furthermore, we collected demonstration data using the proposed teleoperation method. To investigate the effectiveness of utilizing force information in both the input and output of the policy, we conducted imitation learning. As a result, adding force input showed a significant improvement in success rate regardless of the task. On the other hand, the effect of force output depended on the task. Future challenges include improving data collection and optimization methods for parameter identification to more modern methods [36][37], theoretical examination of variable gain control systems, and the investigation of more advanced neural network architectures for imitation learning other than force control systems, as well as compatibility with a wider range of tasks other than those attempted in this study.

### v3

@@ 1 Introduction
However, most of these systems used unilateral control, which only transmits target position values from the leader robot to the follower robot [9]. This method is easy to implement and provides good operability in low-speed, low-contact operations, but it does not provide force feedback and struggles with contact-rich tasks. In fact, it also has difficulties in high-speed operations. On the other hand, there has been an increase in cases where bilateral control, which enables two-way information exchange between the leader robot and the follower robot and provides force feedback, has been adopted. There are various types of bilateral control. In some cases, force reflecting type bilateral control is used, in which external forces from followers are generated feedforward-style on the leader side, like FACTR [10] . However, this method is known to have stability issues, and in many cases, stability is ensured by adding a damping term [10][11], but this tends to impair operability. On the other hand, 4-channel bilateral control is a method in which both the leader and follower transmit their respective positions and forces to each other and perform position control and force control in parallel. Theoretically, a 4-channel architecture can achieve the ideal response of synchronization of position and force perfectly [12][13]. However, to perform position control and force control with opposite objectives simultaneously in a single robot, making them non-interfering requires accurate dynamics models and observation of position, velocity, and external forces. Because the dynamics of manipulators with six or more degrees of freedom involve many parameters and extremely complex functions, obtaining accurate parameters and performing real-time calculations has long been impractical. As a practical implementation method, an accelerated control system was proposed that combines a simplified dynamics model based on linear models of each joint with disturbance estimation (including external force and model error) using a disturbance observer [14][15] and compensation using the estimated disturbance. This enables high performance even when a certain amount of model error exists [16][17][18][19][20]. This acceleration control-based 4-channel bilateral control system has also been studied for application to imitation learning. Sakaino et al. trained a Long-Short-Term Memory (LSTM) [21] using data collected with four-channel bilateral control and conducted experiments to perform contact-rich tasks [22][23][24]. Kobayashi et al. published Alpha-$\alpha$ and Bi-ACT series [25][26][27][28] that are a combination of a low-cost 4-channel bilateral control system Alpha-$\alpha$ and the method to train Action Chunking with Transformer (ACT) [5] using data collected by 4-channel bilateral control.
@@ 2 Methodology
where $\dot{\bm{\theta}}_{-}$ and $\dot{\bm{\theta}}_{+}$ are difference and average of the joint angular velocity. This transformation allows us to think of position and force control independently. In other words, if the same amount of velocity is applied to both the leader and the follower in opposite directions, only the difference in the joint angle will change, not the average. Conversely, if the same amount of velocity is applied to both the leader and the follower in the same direction, only the average joint angle will change, not the difference. Ideally, these do not influence each other, i.e., they are orthogonal, so control systems can be designed separately for each. Since the coordinate transformation matrix is time-invariant, the coordinate transformation of the angle and angular acceleration can be described by the same matrix $\bm{J}_{+-}$. The dynamics of the difference and the average of the joint angles are as follows.
@@
$\begin{bmatrix}\bm{\tau}_{-}\\ \bm{\tau}_{+}\\ \end{bmatrix}_{ext}=\bm{J}_{+-}^{-\top}\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}=\begin{bmatrix}\frac{1}{2}\bm{I}&-\frac{1}{2}\bm{I}\\ \bm{I}&\bm{I}\\ \end{bmatrix}\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}=\begin{bmatrix}\bm{M}_{-}&\bm{0}\\ \bm{0}&\bm{M}_{+}\\ \end{bmatrix}_{d}\begin{bmatrix}\ddot{\bm{\theta}}_{-}\\ \ddot{\bm{\theta}}_{+}\\ \end{bmatrix}+\begin{bmatrix}\bm{D}_{-}&\bm{0}\\ \bm{0}&\bm{0}\\ \end{bmatrix}_{d}\begin{bmatrix}\dot{\bm{\theta}}_{-}\\ \dot{\bm{\theta}}_{+}\\ \end{bmatrix}+\begin{bmatrix}\bm{K}_{-}&\bm{0}\\ \bm{0}&\bm{0}\\ \end{bmatrix}_{d}\begin{bmatrix}\bm{\theta}_{-}\\ \bm{\theta}_{+}\\ \end{bmatrix}$ (6)
where $\bigcirc_{d}$ are desired dynamics parameters, and $\begin{bmatrix}\bm{\tau}_{-},\bm{\tau}_{+}\end{bmatrix}^{\top}_{ext}$ are external forces corresponding to the difference and average of the joint angle. Since the system has position control and force control on orthogonal axes, the system can also be interpreted as position/force hybrid control [32] on transformed coordinates [18]. The control law can be derived by substituting the desired dynamics into the transformed dynamics:
$\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{u}$ $=\begin{bmatrix}\bm{M}_{l}(\bm{\theta}_{l})&\bm{0}\\ \bm{0}&\bm{M}_{f}(\bm{\theta}_{f})\\ \end{bmatrix}\begin{bmatrix}\frac{1}{2}\bm{I}&\bm{I}\\ -\frac{1}{2}\bm{I}&\bm{I}\\ \end{bmatrix}\begin{bmatrix}-\bm{M}_{-d}^{-1}\bm{D}_{-d}\dot{\bm{\theta}}_{-}-\bm{M}_{-d}^{-1}\bm{K}_{-d}\bm{\theta}_{-}\\ \bm{M}_{+d}^{-1}\bm{\tau}_{+}\\ \end{bmatrix}-\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}$
@@
where $\bm{K}_{p}=\frac{1}{2}\bm{M}_{-d}^{-1}\bm{K}_{-d}$, $\bm{K}_{d}=\frac{1}{2}\bm{M}_{-d}^{-1}\bm{D}_{-d}$, and $\bm{K}_{f}=\bm{M}_{+d}^{-1}$. Finally, the control law for each manipulator can be as follows:
$\bm{\tau}_{ref}$ $=\tilde{\bm{M}}(\bm{\theta})\left\{\bm{K}_{p}(\bm{\theta}_{d}-\bm{\theta})+\bm{K}_{d}(\dot{\bm{\theta}}_{d}-\hat{\dot{\bm{\theta}}})+\bm{K}_{f}(\bm{\tau}_{d}+\hat{\bm{\tau}}_{ext})\right\}-\hat{\bm{\tau}}_{ext}+\tilde{\bm{C}}(\bm{\theta},\hat{\dot{\bm{\theta}}})\hat{\dot{\bm{\theta}}}+\tilde{\bm{D}}\hat{\dot{\bm{\theta}}}+\tilde{\bm{g}}(\bm{\theta})$ (8)
@@ 2.2 Velocity and External Force Observer
damping coefficient Here, the observer gain is set using cut-off angular frequency $\omega_{c}$ and the damping coefficient $\zeta$. Converting this to a transfer function expression using the Laplace transform, we can write it as follows.
@@
The critical point is that inertia is not a constant but a function of the joint angle vector. Conventional disturbance observers use constants for inertia, and the effects of modeling errors are included in the estimated external force. Therefore, it was assumed that inertia could be nominalized by feedback compensation of the estimated external force. However, the extent to which modeling errors can be compensated depends on how high the cut-off frequency (observer gain) can be set. In practice, when using high-precision rotary encoders and high sampling frequencies, it is possible to set a fairly high cut-off frequency, and it is practically feasible to ignore modeling errors. However, in the case of low-cost hardware like the one used here, a high cut-off frequency cannot be set, so the error in inertia has a significant impact that cannot be ignored. Pre-identifying the nonlinear dynamics of the manipulator with 38 parameters, including the inertial matrix, realizes high control performance while maintaining a low cost.
Furthermore, using the identified inertia justifies the use of a velocity observer. Although many studies of 4-channel bilateral control used pseudo-differential ($\hat{\dot{\bm{\theta}}}=\frac{\omega_{c}}{s+\omega_{c}}s\bm{\theta}$) for velocity estimation, pseudo-differential is known to cause estimation delays [33]. Nevertheless, pseudo-differential is widely used because the inertia is not accurately known, and methods that can calculate using only angle response values without using inertia are preferred. If accurate inertia can be obtained, it becomes possible to use not only the angular response values but also predictions from the torque reference values, thereby improving the delay in velocity estimation. Improving the delay allows the higher D gain of the position PD control, and setting the D gain higher enables the P gain to be set higher as well, leading to a significant improvement in position control performance.
@@ 3 Experiments
Figure 3: Results of teleoperation comparison. We shook the joint that is closest to the base at high speed by hand. The sign of the external force on the leader is reversed.
Table 1: Mean Absolute Error (MAE) of joint angle, angular velocity, and torque. The proposed method achieved the lowest angle error and the second-lowest angular velocity error. The higher velocity error of the proposed method compared to the symmetric position type can be attributed to the noise in the estimated torque. Note that force errors are natural in free motion. The resolution of the rotary encoder is approximately 0.088 degrees (12-bit, 4096 positions/revolution).
Teleoperation method Angle [deg] Angular velocity [deg/s] Torque [Nm]
Unilateral 2.39 13.1 0.48
Symmetric position type 1.58 7.1 0.50
Force feedback type 1.69 12.5 0.32
4ch using fixed inertia 2.29 28.3 0.26
4ch w/o centrifugal & coriolis comp 0.71 10.1 0.91
4ch using pseudo differential 0.81 48.1 0.61
4ch (proposed) 0.61   9.2 0.52
@@ 3.1 Teleoperation Performance
3.2 Imitation Learning
We experimented to collect demonstration data using the proposed 4-channel bilateral teleoperation method and perform imitation learning. The results are shown in Table 2. We used Action Chunking with Transformer (ACT) [5][25] for neural network architecture. We compare the cases of not using force, using force only for input, and using force for both input and output. Note that since the control system differs between data collection and autonomous operation, a decrease in success rate was expected when force information is not used in the policy output. However, differences in control systems or dynamics between data collection and autonomous operation are not uncommon when using data collection methods other than teleoperation, especially in imitation learning. Additionally, since the dataset’s quality can significantly affect imitation learning results, it is preferable to compare using the same dataset. Considering these points, we conducted this comparison.
Task1: Dual-arm Pick-and-Place. We performed a task of grasping blocks of different widths between two arms. We prepared five types of blocks, ranging from 10 mm to 50 mm in width, and collected a total of 10 demonstration data sets, two for each type. The blocks are shown in Fig. 8. As a result, when no force information was used for either input or output, picking was not possible in cases with small widths of 10 mm and 20 mm. In cases with relatively large widths of 30 mm to 50 mm, picking was possible once, but problems such as dropping the object immediately or the angle changing while holding it occurred, and the object could not be placed in the correct position. When force information was used in the input, the success rate of picking improved significantly, with all trials resulting in successful picking. However, instability after picking persisted, and placement failures occurred frequently in cases where the width was farthest from the average of 10mm and 50mm. When both input and output used force information, both pick and place were successful in all trials. The stability of holding after picking improved, with no instances of dropping midway or misalignment.
Task2: Nut Turning. We tried the nut turning task as a task that requires more force and speed. As a result, when no force was applied to either the input or output, the experiment was completely unsuccessful, but when force was applied to the input, the experiment was successful regardless of whether force was applied to the output. When no force was applied to the input, the hand often stopped at the position where it overlapped with the nut on the bird’s-eye view image, and it seemed difficult to touch the nut due to insufficient force information.
Task3: Cucumber Peeling. As a task involving handling irregularly shaped objects, we tried peeling cucumbers with a peeler. As a result, when no force information was used for either input or output, the experiment was completely unsuccessful. When force information was used for input, the experiment was successful in 3/5 cases, and when force information was used for both input and output, the experiment was successful in 2/5 cases. The main failure case was when the system transitioned to the finished state before touching the cucumber or peeler. This is thought to be because there was not much difference between the image at the start of the task and the image after the task was completed.
Table 2: Success rate of imitation learning tasks by Action Chunking with Transformer
@@ 3.2 Imitation Learning
block width pick place pick place pick place
Dual-arm Pick-and-Place 10 mm 0/5 0/5 5/5 0/5 5/5 5/5
20 mm 0/5 0/5 5/5 5/5 5/5 5/5
30 mm 5/5 0/5 5/5 5/5 5/5 5/5
40 mm 5/5 0/5 5/5 4/5 5/5 5/5
50 mm 5/5 0/5 5/5 0/5 5/5 5/5
@@ 4 Conclusion
In this study, we demonstrated that a low-cost manipulator can achieve high-speed teleoperation with force feedback without force sensors by incorporating a high-precision manipulator dynamics model into a four-channel bilateral teleoperation system. We compared the proposed method to three teleoperation methods, such as unilateral control, symmetric position type, and force feedback type, and a 4-channel bilateral control without identified manipulator dynamics. We confirmed that the proposed method demonstrates higher position tracking performance during high-speed operations. Furthermore, we collected demonstration data using the proposed teleoperation method. To investigate the effectiveness of force information in both the input and output of the policy, we conducted imitation learning. As a result, adding force input showed a significant improvement in success rate regardless of the task. On the other hand, the effect of force output depended on the task. Future challenges include improving data collection and optimization methods for parameter identification to more modern approaches [36][37], and applying more advanced neural network architectures for imitation learning, as well as enhancing compatibility with a broader range of tasks beyond those explored in this study.